You wish to conduct a hypothesis test to determine if a bivariate data set has a significant correlation among the two variables. That is, you wish to test the claim that there is a correlation (Ha:ρ≠0Ha:ρ≠0). You have a data set with 69 subjects, in which two variables were collected for each subject. You will conduct the test at a significance level of α=0.05α=0.05.

Find the critical value for this test (Pick the closest conservative value from the table).
rc.v. = ±±

Report answers accurate to three decimal places.
Answer #1

The sample size is n=69, so then the number of degrees of freedom is df=n−2=69−2=67

The corresponding critical correlation value rc​ for a significance level of α=0.05, for a two-tailed test is:

rc​=0.237

Observe that in this case, the null hypothesis is rejected if ∣r∣>rc​=0.237.